OpenStudy (anonymous):

How the hell do you calculate the missing sides of a pyramid/cone if you just have the volume? Please help. Math will be the death of me. For example, a cylinder when the height is 3 and the Volume is 339.3 ... I'm supposed to be finding the radius.

OpenStudy (anonymous):

For a cylinder: Volume = pi*r^2*h. Solving for r would get: r = (V/(pi*h))^(1/2) since you know h and V, you can find r. In your example, (339.3/(3pi))^(1/2) = (36)^(1/2) = 6

OpenStudy (anonymous):

THANK YOU. I LOVE YOU.

OpenStudy (anonymous):

Wait, I don't understand why you would go ^(1/2), but I get everything up to that point.

OpenStudy (anonymous):

^(1/2) is the same as taking the square root of the entire thing, since you did have an r^2 to begin with. To get just r, you must take the square root of both sides.

OpenStudy (anonymous):

Ohh okay, thank you